OpenStudy (anonymous):

Find the slope and y-intercept of the line in the given equation. y=-/10x-1

OpenStudy (anonymous):

yint=-1 but ur slope is weird it has division after - sign

OpenStudy (anonymous):

\[y=-1/10x-1\]

OpenStudy (anonymous):

no it is over1/10

OpenStudy (anonymous):

y = mx + b m = gradient (slope), b = y-intercept In your case, gradient = -1/10, y-intercept = -1